# Taxonomizing Representational Harms using Speech Act Theory

### v1

@@ 3 A Granular Taxonomy of Stereotyping, Demeaning, and Erasure
Threatens or promises exclusion "You [social group] better follow these rules or you’ll be kicked out"
@@
Denies fair access (excludes) "No [social group] allowed"/"[Social group] only"
@@
Table 1: Our taxonomy of stereotyping, demeaning, and erasure as types of illocutionary acts. We further divide each type into illocutionary act patterns that share illocutionary effects (as well as the shared perlocutionary effect of entrenching of one or more harmful social hierarchies). We provide an example for each illocutionary act pattern.
@@
Following Searle and Vanderveken (1985), who further divided the five basic classes of illocutionary acts into subclasses, [^7] Searle further divided representatives, expressives, directives, commissives, and declarations into a set of granular subclasses (Searle and Vanderveken, 1985). For example, commissives include promises, threats, vows, etc., while directives include commands, demands, warnings, etc. This taxonomical structure parallels those of grammatical categories and linguistic forms. For more information about these subclasses, see the work of Searle and Vanderveken (1985). we draw on the definitions in the previous section to analyze stereotyping, demeaning, and erasure from multiple altitudes. Specifically, we divide the three types of illocutionary acts described in the previous section—stereotyping, demeaning, and erasure—into subtypes, namely illocutionary act patterns. These patterns highlight distinguishing features of the types of illocutionary acts, reflecting the specific ways they arise via language. Together, these levels capture an utterance’s purpose, while differing in their conceptual altitudes. Crucially, they provide the granularity needed to develop valid measurement instruments and effective mitigations by supporting the identification of relevant utterances that vary in terms of other salient dimensions, such as social hierarchies, social groups, and characteristics.
For example, the utterance we used to illustrate our definition of stereotyping in Section 2.2.1—i.e., “Gay men are not very masculine”— is a representative (and stereotyping) illocutionary act, whose illocutionary act pattern is describing a social group stereotypically. Similarly, the utterance we used to illustrate our definition of demeaning—i.e., “Obviously bisexuals are messed up”—is a representative (and demeaning) illocutionary act that stigmatizes a social group on the basis of deviance. The utterance we used to illustrate our definition of erasure in Section 2.2.3—i.e., “There is no way that bisexuality is a real thing”—is a representative (and erasing) illocutionary act that denies the existence of a social group. Together, these examples illustrate the relationships between illocutionary acts and illocutionary act patterns.
@@ 4.1 Case Study: Conceptual Debates
In this case study, we use our framework and taxonomy to analyze three taxonomies of “representational harms” [^9] We use quotation marks to indicate that others’ conceptualizations of representational harms may differ from ours. presented in previous work (Chien and Danks, 2024; Blodgett, 2021; Katzman et al., 2023), paying particular attention to those that focus on system behaviors and are therefore most similar to our taxonomy. We provide visual summaries of the three taxonomies, which we selected for their coverage of system behaviors and real-world impacts, in Appendix G. Analyzing these taxonomies highlights how our framework and taxonomy can support the identification and development of granular taxonomies that possess internal theoretical coherence by 1) drawing clearer distinctions between system behaviors and their impacts and 2) providing relevant materials for identifying different levels of granularity within taxonomies of system behaviors.
@@
Katzman et al. (2023) Stereotyping, demeaning, erasure Reifying social groups, denying people the opportunity to self-identify
@@ 4.1.1 A Taxonomy of Real-World Impacts
Chien and Danks (2024) presented a taxonomy of individual, interpersonal, and social harms caused by representations of social groups, arguing in favor of conceptualizing representational harms as the (negative) real-world impacts on people’s psychological states by particular types of system behaviors. They also highlighted the need to align mitigations with a comprehensive understanding of these impacts. According to our framework, their taxonomy focuses exclusively on perlocutionary effects. Structured as a granular, multilevel taxonomy with three top-level types of impacts—i.e., those affecting people’s understandings of identity, those affecting people’s stress levels due to perceived danger or lack of control, and those affecting people’s feelings about their own identities and interpersonal relationships within and between social groups—these types are further divided into more granular patterns. The taxonomy therefore complements our taxonomy by focusing on a different object of study, namely perlocutionary effects.
@@ 4.1.3 Summary
By using our framework and taxonomy to analyze three existing taxonomies of “representational harms,” we shed light on the ways in which these taxonomies differ from one another (see Table 2 for a summary of these differences) and from our taxonomy in terms of their objects of study and granularities. In this way, our framework and taxonomy provide a theoretical basis for debates about what constitutes a representational harm and how such harms should be measured, laying the groundwork for the operationalization process.
@@ 5 Conclusion
Although representational harms are widely recognized among fairness-related harms caused by generative language systems, definitions of these harms are commonly under-specified and sometimes conflate system behaviors and their impacts. This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. To address this challenge, we presented a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms as the perlocutionary effects, (i,e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). We then used this framework to develop a granular taxonomy of illocutionary acts that highlights distinguishing features of stereotyping, demeaning, and erasure, going beyond the high-level taxonomies presented in previous work. We discussed the ways that our framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. To summarize, measuring and mitigating representational harms requires granular taxonomies that possess internal theoretical coherence. Speech act theory provides a coherent conceptual framework for identifying and developing such taxonomies.
@@ 6 Limitations
Our framework, grounded in speech act theory Austin (1962), conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). However, generative language systems often support myriad use cases and generative AI systems often incorporate other modalities, such as speech and vision. We focus specifically on language generation systems, without accounting for modalities beyond language. Our conceptualization of representational harms also focuses on the occurrence of particular system behaviors, rather than taking a distribution- or performance-based approach (Katzman et al., 2023). We believe our framework applies across these dimensions of variation, but would require adaptation to focus on propositions instead of utterances. We defer a detailed exploration of this direction to future work.
@@
In Section 4, we discussed the ways that our framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. Such conceptual debates lay the groundwork for the operationalization process by providing a clear specification of exactly what should be operationalized and why. Although we chose to focus on conceptual debates, the operational debates are also crucial to developing valid measurement instruments. We therefore hope that future work picks up from where we left off, focusing on the role of our framework and taxonomy in this part of the measurement process.
Finally, we emphasize that our collective positionalities influence our thinking about language as a social phenomenon, representational harms, and language generation systems. We are an interdisciplinary group of linguists with a variety of specializations; experts in natural language processing and machine learning; and applied scientists. As a result, our shared perspective is broad and has been influenced by our myriad conversations with other technologists, computational social scientists, linguists, natural language processing and machine learning experts, applied scientists, and engineers. From a linguistic perspective, we are oriented toward functional approaches to language variation, pragmatics, linguistic anthropology, sociolinguistics, and philosophy, as well as the broader social sciences and humanities. That said, despite our broad shared perspective, there are likely gaps in our thinking due to our collective positionalities.
@@ 7 Acknowledgments
We thank Solon Barocas, Su Lin Blodgett, Tim Vieira, Philip Resnik and his students, and many others who have shaped this paper over the years.
@@ A.1 Components of Speech Act Theory
An utterance’s illocutionary act corresponds to its purpose, often characterized by what its speaker intends [^15] We emphasize that generative language systems are not sentient and do not have intent. We address this tension between the nature of generative language systems and intent-centric understandings of speech act theory in Appendix A.2. to accomplish through the production of that utterance.161616 Because humans cannot read minds, the illocutionary act is often the site of significant linguistic and social ambiguity. Interpreting the illocutionary act can therefore require the hearer to make inferences based on word choice and ordering, other linguistic and paralinguistic cues, and social context. Building on Austin’s presentation of speech acts, Searle (1976) identified and described five basic classes of illocutionary acts: representatives, expressives, directives, commissives, declarations. For example, the illocutionary act for the utterance “Can you pass the salt?” is a directive—i.e., a request for an action to be carried out—namely that the salt be passed. Because an illocutionary act can take many different forms, it can therefore correspond to many different locutionary acts. For example, if a speaker wants the hearer to pass the salt, they might say “Can you pass the salt?” but they might instead say “This food needs salt!”, “Can you give me that?”, “May I have the salt, please?”, “Give me the salt now!”, “Salt! Immediately!”, and so on.
@@ A.2 A Speech Act Theoretical Model of Communication
Speech act theory provides the necessary scaffolding for the development of a basic model of interactive communication between humans, which includes the the word choice and ordering in a given utterance, the purpose of the utterance, and the real-world impacts of the utterance. This model is helpful in understanding how utterances—including those that cause representational harms—convey meaning between speakers and hearers. We argue that parts of this model can be mapped to interactions between systems and humans. Below, we explain this model of communication in human-to-human interactions and then map it onto system-to-human interactions, where the system is considered the “speaker” and the human user is the “hearer”. The model is shown in Figure 2.
Figure 2: An interactive model of communication based in speech act theory.
@@
However, mapping human-to-human interaction to system-to-human interaction requires careful consideration of the role of speaker intention, which is commonly integrated into conceptualizations of illocution. While some traditional schools of thought within speech act theory theorize the meaning of an utterance as the intention of the speaker—i.e., within illocution (see Searle and Vanderveken, 1985; Grice, 1957, 1968, 1969; Sperber and Wilson, 1995; Neale, 1992; Stalnaker, 1999, inter alia)—these theories of meaning cannot be used in contexts where the “speaker” is a generative language system without inadvertently ascribing intent to these systems. We, like others, have taken issue with ascribing intention to generative language systems (Bender and Koller, 2020; Bender et al., 2021) and thus believe these theories of meaning are inappropriate when the “speaker” is non-sentient.
@@
The model of communication presented here is functional for human-to-human and system-to-human communication, centering interpretation of meaning as uptake by the human hearer. For the human hearer, the expression, or utterance, is the locution, while its purpose—comprised of the inferred meaning and inferred communicative intent, and the entailed consequences of those inferences—reflect the illocution, and the actual real-world impacts of the utterance are the perlocutionary effects.
While not treated in detail here, this communicative model may be adapted to other generative modalities (e.g., image, speech, video). For example, visual representations that are outputs of image generation systems can be conceptualized as types of illocutionary acts whose meaning is interpreted by a human “hearer” or recipient and that can have any number of perlocutionary effects.
@@ A.3 A Speech Act Theoretical Perspective on Conversational Ideals
Representatives commit the speaker to the belief that the stated representation corresponds to some true state of affairs in the world, thus communicating observations and beliefs. Expressives are illocutionary acts that express a psychological or subjective state, as in expressions of joy, frustration, or sadness. In contrast to representatives and expressives that present some perspective on the world, directives are illocutionary acts that direct or instruct the listener to take some action. These are common in interactions involving advice or how-to instructions, but the class of directives also includes illocutionary acts that prohibit or forbid the hearer from taking some action. Commissives are distinct in that they commit the speaker to a future action—e.g., promises and threats. Finally, declarations—sometimes called performatives—change or update some aspect of world to match some part of the utterance, such as in ceremonial pronouncements like baptisms, ship namings, and weddings that change the legal status of some entity or set of entities.
Recently, many have acknowledged the relevance of pragmatics in the design and evaluation of generative language systems (Pautler and Quilici, 1998; Goodman and Frank, 2016; Freiman and Miller, 2019; Sap et al., 2020; Fried et al., 2023; Kasirzadeh and Gabriel, 2023). Of these, Kasirzadeh and Gabriel (2023) draw special attention to speech act theory by highlighting the kinds of illocutionary acts—namely, the classes identified by Searle (1976)—that are inappropriate when produced by generative language systems due to their lack of embodiment and psychology. Specifically, they argue that three of Searle’s classes should not be present in the outputs of generative language systems: expressives—due to generative language systems’ lack of internal psychology; commissives—due to generative language systems’ inability to follow through beyond episodic interactions; and declarations—due to generative language systems’ lack of authority to bring about (specifically, ceremonial or legal) changes in the world. According to this view, different kinds of speech acts reflect different communicative goals, and some of these goals cannot be achieved by generative language systems precisely because they are not human.

### v2

@@ 3 A Granular Taxonomy of Stereotyping, Demeaning, and Erasure
Threatens or promises exclusion “You [social group] better follow these rules or you’ll be kicked out”
@@
Denies fair access (excludes) “No [social group] allowed” / “[Social group] only”
@@
Table 1: Our taxonomy of stereotyping, demeaning, and erasure as types of illocutionary acts. We further divide each type into illocutionary act patterns that share illocutionary effects (as well as the shared perlocutionary effect of entrenching one or more harmful social hierarchies). We provide an example for each illocutionary act pattern.
@@
Following Searle and Vanderveken (1985), who further divided the five basic classes of illocutionary acts into subclasses, [^8] Searle further divided representatives, expressives, directives, commissives, and declarations into a set of granular subclasses (Searle and Vanderveken, 1985). For example, commissives include promises, threats, vows, etc., while directives include commands, demands, warnings, etc. This taxonomical structure parallels those of grammatical categories and linguistic forms. For more information about these subclasses, see the work of Searle and Vanderveken (1985). we draw on the definitions in the previous section to analyze stereotyping, demeaning, and erasure from multiple altitudes. Specifically, we divide the three types of illocutionary acts described in the previous section—stereotyping, demeaning, and erasure—into subtypes, namely illocutionary act patterns. These patterns highlight distinguishing aspects of the types of illocutionary acts, reflecting the ways they manifest in language. Together, these levels capture an utterance’s purpose, while differing in their conceptual altitudes. Crucially, they provide the granularity needed to develop valid measurement instruments and effective mitigations by supporting the identification of relevant utterances that vary in terms of other salient dimensions, such as social hierarchies, social groups, and characteristics.
For example, the utterance we used to illustrate our definition of stereotyping in Section 2.2.1—i.e., “Gay men are not very masculine”— is a representative (and stereotyping) illocutionary act, whose illocutionary act pattern is describing a social group stereotypically, or in an essentializing way. Similarly, the utterance we used to illustrate our definition of demeaning—i.e., “Obviously bisexuals are messed up”—is a representative (and demeaning) illocutionary act that stigmatizes a social group on the basis of deviance. The utterance we used to illustrate our definition of erasure in Section 2.2.3—i.e., “There is no way that bisexuality is a real thing”—is a representative (and erasing) illocutionary act that denies the existence of a social group. Together, these examples illustrate the relationships between illocutionary acts and illocutionary act patterns.
@@ 4.1 Case Study: Conceptual Debates
In this case study, we use our framework and taxonomy to analyze three taxonomies of “representational harms” [^10] We use quotation marks to indicate that others’ conceptualizations of representational harms may differ from ours. presented in previous work (Chien and Danks, 2024; Blodgett, 2021; Katzman et al., 2023), paying particular attention to those that focus on system behaviors and are therefore most similar to our taxonomy. We provide visual summaries of the three taxonomies, which we selected for their coverage of system behaviors and real-world impacts, in Appendix I. Analyzing these taxonomies highlights how our framework and taxonomy can support the identification and development of granular taxonomies that possess internal theoretical coherence by 1) drawing clearer distinctions between system behaviors and their impacts and 2) providing relevant materials for identifying different levels of granularity within taxonomies of system behaviors.
@@
Katzman et al. (2023) Stereotyping, demeaning, erasure Reifying social groups, denying people the opportunity to self-identify n/a
@@ 4.1.1 A Taxonomy of Real-World Impacts
Chien and Danks (2024) presented a taxonomy of individual, interpersonal, and social harms caused by representations of social groups, arguing in favor of conceptualizing representational harms as the (negative) real-world impacts of particular types of system behaviors on people’s psychological states, including cognitive, affective, and emotional states. They also highlighted the need to align mitigations with a comprehensive understanding of these impacts. According to our framework, their taxonomy focuses exclusively on perlocutionary effects. Structured as a granular, multilevel taxonomy with three top-level types of impacts—i.e., those affecting people’s understandings of identity, those affecting people’s stress levels due to perceived danger or lack of control, and those affecting people’s feelings about their own identities and interpersonal relationships within and between social groups—these types are further divided into more granular patterns. The taxonomy therefore complements our taxonomy by focusing on a different object of study, namely perlocutionary effects.
@@ 4.1.3 Summary
By using our framework and taxonomy to analyze three existing taxonomies of “representational harms,” we shed light on the ways in which these taxonomies differ from one another (see Table 2 for a summary of these differences) and from our taxonomy in terms of their objects of study and their granularities. In this way, our framework and taxonomy provide a theoretical basis for debates about what constitutes a representational harm and how such harms should be measured, laying the groundwork for engaging with existing measurement instruments and developing new ones. In Appendix G, we use our framework and taxonomy to analyze one such existing instrument for measuring stereotyping and demeaning—the FairPrism dataset (Fleisig et al., 2023)—and its underlying definitions. In Appendix H, we outline a series of decisions to make and steps to take to turn our framework and taxonomy into new measurement instrument—namely, a set of annotation guidelines.
@@ 5 Conclusion
Although representational harms are widely recognized among fairness-related harms caused by generative language systems, definitions of these harms are commonly under-specified and sometimes conflate system behaviors and their impacts. This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. To address this challenge, we made a theoretical contribution to the specification of representational harms by introducing a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms as the perlocutionary effects, (i,e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). We then used this framework to develop a granular taxonomy of illocutionary acts that highlights distinguishing aspects of stereotyping, demeaning, and erasure, going beyond the high-level taxonomies presented in previous work. We discussed the ways that our theory-grounded framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. To summarize, measuring and mitigating representational harms requires granular taxonomies that possess internal theoretical coherence. Speech act theory provides a coherent conceptual foundation for identifying and developing such taxonomies.
@@ 6 Limitations
Our framework, grounded in speech act theory Austin (1962), conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). However, generative language systems often support myriad use cases and generative AI systems often incorporate other modalities, such as speech and vision. We focus specifically on language generation systems, without accounting for modalities beyond language. Our conceptualization of representational harms also focuses on the occurrence of particular system behaviors, rather than taking a distribution- or performance-based approach (Katzman et al., 2023). We believe our framework applies across these dimensions of variation, but it would require some adaptation to focus on propositions instead of utterances. We defer a detailed exploration of this direction to future work.
@@
In Section 4, we discussed the ways that our framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. Such conceptual debates lay the groundwork for the operationalization process by providing a clear specification of exactly what should be operationalized and why. Although we chose to highlight our theoretical contribution by focusing on conceptual debates, operational debates are also crucial to developing valid measurement instruments. We offer in Appendix G and Appendix H insight into the ways our framework and taxonomy can be used to analyze existing measurement instruments and develop new ones. We hope that future work picks up from where we left off, focusing on the role of our framework and taxonomy in this part of the measurement process.
Finally, we emphasize that our collective positionalities influence our thinking about language as a social phenomenon, representational harms, and language generation systems. We are an interdisciplinary group of linguists with a variety of specializations, experts in natural language processing and machine learning, and applied scientists. As a result, our shared perspective is broad and has been influenced by our myriad conversations with other technologists, computational social scientists, linguists, natural language processing and machine learning experts, applied scientists, and engineers. From a linguistic perspective, we are oriented toward functional approaches to language variation, pragmatics, linguistic anthropology, sociolinguistics, and philosophy, as well as the broader social sciences and humanities. That said, despite our broad shared perspective, there are likely gaps in our thinking due to our collective positionalities.
@@ 7 Acknowledgments
We thank Solon Barocas, Su Lin Blodgett, Tim Vieira, Philip Resnik and his students, and many others who have shaped this paper over the years. We also thank our reviewers for their feedback.
@@ A.1 Components of Speech Act Theory
An utterance’s illocutionary act corresponds to its purpose, often characterized by what its speaker intends [^16] We emphasize that generative language systems are not sentient and do not have intent. We address this tension between the nature of generative language systems and intent-centric understandings of speech act theory in Appendix A.2. to accomplish through the production of that utterance.171717 Because humans cannot read minds, the illocutionary act is often the site of significant linguistic and social ambiguity. Interpreting the illocutionary act can therefore require the hearer to make inferences based on word choice and ordering, other linguistic and paralinguistic cues, and social context. Building on Austin’s presentation of speech acts, Searle (1976) identified and described five basic classes of illocutionary acts: representatives, expressives, directives, commissives, declarations. For example, the illocutionary act for the utterance “Can you pass the salt?” is a directive—i.e., a request for an action to be carried out—namely that the salt be passed. Because an illocutionary act can take many different forms, it can therefore correspond to many different locutionary acts. For example, if a speaker wants the hearer to pass the salt, they might say “Can you pass the salt?” but they might instead say “This food needs salt!”, “Can you give me that?”, “May I have the salt, please?”, “Give me the salt now!”, “Salt! Immediately!”, and so on.
@@ A.2 A Speech Act Theoretical Model of Communication
Speech act theory provides the necessary scaffolding for the development of a basic model of interactive communication between humans, which includes the word choice and ordering in a given utterance, the purpose of the utterance, and the real-world impacts of the utterance. This model is helpful in understanding how utterances—including those that cause representational harms—convey meaning between speakers and hearers. We argue that parts of this model can be mapped to interactions between generative language systems and humans. Below, we explain this model of communication in human-to-human interactions and then map it onto system-to-human interactions, where the system is considered the “speaker” and the human user is the “hearer”. The model is illustrated below in Figure 2.
Figure 2: An interactive model of communication based on speech act theory.
@@
However, mapping human-to-human interaction to system-to-human interaction requires careful consideration of the role of speaker intention, which is commonly integrated into conceptualizations of illocution. While some traditional schools of thought within speech act theory theorize the meaning of an utterance as the intention of the speaker—i.e., within illocution (see Searle and Vanderveken, 1985; Grice, 1957, 1968, 1969; Sperber and Wilson, 1995; Neale, 1992; Stalnaker, 1999, inter alia)—these theories of meaning cannot be used in contexts where the “speaker” is a generative language system without inadvertently ascribing intent to such systems. We, like others, take issue with ascribing intention to generative language systems (Bender and Koller, 2020; Bender et al., 2021) and thus believe these theories of meaning are inappropriate when the “speaker” is non-sentient.
@@
The model of communication presented here is functional for human-to-human and system-to-human communication, centering interpretation of meaning as uptake by the human hearer. For the human hearer, the expression, or utterance, is the locution, while its purpose—comprised of the inferred meaning and inferred communicative intent, and the entailed consequences of those inferences—reflect the illocution. The actual real-world impacts of the utterance are the perlocutionary effects.
While not treated in detail here, this communicative model may be adapted to other generative modalities (e.g., image, speech, video). For example, visual representations that are outputs of image generation systems can be conceptualized as types of illocutionary acts whose meaning is interpreted by a human “hearer” or recipient, resulting in any number of perlocutionary effects.
@@ A.3 A Speech Act Theoretical Perspective on Conversational Ideals
Representatives commit the speaker to the belief that the stated representation corresponds to some state of affairs in the world, thus communicating observations and beliefs. Expressives are illocutionary acts that express a psychological or subjective state, as in expressions of joy, frustration, or sadness. In contrast to representatives and expressives that present some perspective on the world, directives are illocutionary acts that direct or instruct the listener to take some action. These are common in interactions involving advice or how-to instructions. Notably this class of directives also includes illocutionary acts that prohibit or forbid the hearer from taking some action. Commissives are distinct in that they commit the speaker to a future action—e.g., promises and threats. Finally, declarations—sometimes called performatives—change or update some aspect of world to match some part of the utterance, such as in ceremonial pronouncements like baptisms, ship namings, and weddings that change the legal status of some entity or set of entities.
Recently, many have acknowledged the relevance of pragmatics in the design and evaluation of generative language systems (Pautler and Quilici, 1998; Goodman and Frank, 2016; Freiman and Miller, 2019; Sap et al., 2020; Fried et al., 2023; Kasirzadeh and Gabriel, 2023). Of these, Kasirzadeh and Gabriel (2023) draw special attention to speech act theory by highlighting the kinds of illocutionary acts—namely, the classes identified by Searle (1976)—that are inappropriate when produced by generative language systems due to their lack of embodiment and psychology. Specifically, they argue that three of Searle’s classes should not be present in the outputs of generative language systems: expressives—due to generative language systems’ lack of internal psychology; commissives—due to generative language systems’ inability to follow through beyond a given interaction; and declarations—due to generative language systems’ lack of authority to bring about (specifically, ceremonial or legal) changes in the world. According to this view, different kinds of speech acts reflect different communicative goals, and some of these goals cannot be achieved by generative language systems precisely because they are not human.